OpenStudy (anonymous):

I need help...find the general term a, for the geometric sequence. a{1}=-3 and a{2}=15

OpenStudy (anonymous):

geometric means each term is multiplied by the same number \(r\) to get the next term since \(a_2=a_1r\) you can find \(r\) via \(r=\frac{a_2}{a_1}\)

OpenStudy (anonymous):

in your case \(r=\frac{15}{-3}=-5\) so general term looks like \[a_n=-3\times (-5)^{n-1}\]

OpenStudy (anonymous):

oh so I basically divide the numbers? and thats the general term?? so confusing.